is irt possible that 4×10-19cb charge is present on any body ?

Grade:12

3 Answers

Saurabh Koranglekar
3 years ago
Dear student

No it is not possible as it should be in multiple of 1.6 * 10^(-19)

Regards
Arun
3 years ago
Dear student
 
As we know
 
Q = ne
 
4* 10^-19 = n * 1.6 * 10^-19
 
n = 4/1.6
 
As this is not an integer. Hence this cam not be possible
 
Hope it helps
Vikas TU
3 years ago
Dear student 
No, as the minimum possible is 1.6 x 10^-19, which is the charge of an electron. a body cant have charge without the integral multiples of the charge of elec. As, n is in fraction which is not possible. Hence, a body can''t have a charge of 0.8X10^(-19).
Good Luck